a ball of mass 200g is moving with velocity 10m/s if error in measurement of velocity is 0.1% the uncertainity in position in metres is ?

Grade:11

4 Answers

Naveen Kumar
9 years ago
Mass of Ball (m)= 0.2kg
Velocity of the Ball=10m/s
ΔV=0.1%=10*0.1/100=0.01m/s
Hence, uncertainity in momentum,Δp=mΔV= 0.2*0.01kg-ms-1=0.002kg-ms-1
Let change in position=Δx
So by uncertainty principle,Δp*Δx=h/4π
…............................Δx=h/(4π*Δp)
..........................................Δx=6.64*10-34/(4*3.14*0.002)m
KARTIK GUPTA
6 years ago
The above equation is perfect. The answer comes out to be on solving the equation~~(around) :1.3*10^-31
Bijaylaxmi mishra
6 years ago
Mass = 200g = 0.200kg Speed = 20cm = 0.20m/sKinetic energy = 1/2mv^2 = 1/2 (0.200kg )(0.20m/s)^2 = 0.004 joule.
GAWAND VIDHISHA
5 years ago
THE ABOVE EQUATION IS CORRECT BUT THE SOLVING AND CALCULATION PART IS 
6.636 X10-34 / 4X3.14X0.002
= 6636X 10-32/ 2412
= 2.7 X 10-32